### v1

@@ Abstract
In Split Federated Learning (SFL), the clients collaboratively train a model with the help of a server by splitting the model into two parts. Part-1 is trained locally at each client and aggregated by the aggregator at the end of each round. Part-2 is trained at a server that sequentially processes the intermediate activations received from each client. We study the phenomenon of catastrophic forgetting (CF) in SFL in the presence of data heterogeneity. In detail, due to the nature of SFL, local updates of part-1 may drift away from global optima, while part-2 is sensitive to the processing sequence, similar to forgetting in continual learning (CL). Specifically, we observe that the trained model performs better in classes (labels) seen at the end of the sequence. We investigate this phenomenon with emphasis on key aspects of SFL, such as the processing order at the server and the cut layer. Based on our findings, we propose Hydra, a novel mitigation method inspired by multi-head neural networks and adapted for the SFL’s setting. Extensive numerical evaluations show that Hydra outperforms baselines and methods from the literature.
@@ 1 Introduction
Split Learning (SL) is a distributed learning method where part of the clients’ training is offloaded to a server vepakomma2018split; gupta2018distributed. This is particularly useful in cases where clients’ resources are insufficient to perform on-device training. In a nutshell, in SL, the deep neural network (NN) model is split into two parts of consecutive layers: part-1 and part-2, where part-1’s last layer is called the cut layer. These parts are trained by the client and the server, respectively. There is a plethora of proposed variations for SL that mostly differ in the way the training at the server takes place. In this work, we focus on Split Federated Learning (SFL) and, in particular, SplitFedv2 thapa2022splitfed, which is one of the most prevailing versions hafi2024split. In fact, due to its workflow, SFL has been shown to deliver a faster convergence than other variants thapa2022splitfed.
@@
Motivation. Part-1 is trained as in Federated Learning (FL), while part-2 is trained over the activations that clients send. This dual aspect of SFL makes it particularly susceptible to catastrophic forgetting (CF) in the presence of data heterogeneity, a common phenomenon in training NNs goodfellow2013empirical. In practice, part-1 and part-2 may suffer from CF differently, as depicted in Fig. 1 (bottom). Like in FL, the local updates of part-1 drift away from the global optima as a result of the aggregation (e.g., FedAvg mcmahan2017communication) at the end of every round. On the other hand, our experiments reveal a new finding: the processing order at the server has a significant impact on CF in part-2. This CF reminds the one of Continual Learning (CL) wang2024comprehensive: the model parameters perform well for the dominant label of the last processed client, but forget the previously seen labels, i.e., in every update step at the server, part-2 tends to forget the knowledge learned during the previous steps, leading to a performance gap among labels.
@@
Fig. 2 showcases the CF in SFL as a result of the processing order at the server under non-independently and identically distributed (non-IID) data. In particular, we assume that the data of each client contains a highly represented (dominant) label. Each line depicts the accuracy of the labels that are dominant in clients that are processed at position 1 to 10 (in CIFAR-10 with 10 labels), where details on the experimental setup will be given in Sec. 3.1. We see that, under non-IID, the labels that are dominant at the clients who are processed last at the server consistently outperform the other labels. The disparity in performance between the labels seen at the end of the sequence and those seen earlier in the sequence is related to the CF in part-2. We identify this as intra-round catastrophic forgetting (intraCF), as it occurs at a different granularity (i.e., within a round) than the model drifting of part-1, which also contributes to the phenomenon. Moreover, the global accuracy is heavily affected, i.e., $43\%$ after 100 rounds, whereas, under IID, the global accuracy reaches $80\%$ after 100 rounds.
@@
(i) Metrics of forgetting & Mitigation Methods. Related work on forgetting in SL, e.g., feng2023iotsl; xia2024korea focuses solely on the global accuracy, ignoring the disparity (gap) of performance among labels. Further, methods proposed to mitigate forgetting and/or data heterogeneity in CL and FL, such as EWC kirkpatrick2017overcoming and Scaffold karimireddy2020scaffold, cannot be applied in the setting of SFL as the former is based on the notion of temporal data and the latter requires the entire model.
(ii) The roles of the cut layer and processing order. The majority of related work assumes that the server processes the clients’ data in a random order thapa2022splitfed; liao2023mergesfl and often ignores the role of the cut layer in CF. Moreover, existing theoretical analysis on SFL is limited to convergence analysis under standard assumptions without focusing on these key elements of SFL hanconvergence.
(iii) Empirical evaluation. The existing theoretical findings are often showcased in limited evaluation scenarios (e.g., partition methods), see hanconvergence; li2023convergence. Further, it is unclear whether some of the theoretical claims proved in the settings of CL and FL may hold in SFL.
@@
2. Providing valuable insights into this phenomenon (e.g., wrt the impact of cut layer and processing order at the server) and comparing them with existing theoretical findings. To this end, we focus on the accuracy, the performance gap among labels, and the metric of backward transfer. 3. Proposing Hydra, a novel mitigation method that trains multiple copies of the last layers of part-2 after grouping the clients’ input based on their data distributions. Hydra is designed based on the insights from our analysis and induces minimal overhead (which can be further reduced).
@@ 2 Related Work
In this section, we discuss related work on data heterogeneity and CF in FL and SL. Several works study CF in FL under non-IID data as a result of model drifting and client selection policies. Mitigation methods utilize knowledge distillation lee2022preservation; ma2022continual and regularization xu2022acceleration, among others. We notice, however, that most of the methods for FL require control of the entire model, e.g., karimireddy2020scaffold; zhao2018federated; hu2024fedmut, which is incompatible with the SFL workflow. Moreover, optimization techniques in FL wang2020tackling; li2020federated assume a symmetric training workflow, with full local steps, or per-client regularization terms, which also is not applicable for SFL. On the other hand, replay-based methods for FL, e.g., qibetter; li2024towards induce computational and memory overhead. Concerning SL (in its various definitions), CF has been mostly studied in terms of global accuracy under non-IID data. Proposed solutions are based on generated data and knowledge transfer for SplitNN feng2023iotsl; feng2024slwf, as well as on knowledge replay for SplitFedv1 xia2024korea. Further, madaan2022vulnerability provides a preliminary study of the impact of the processing order in different versions of SL. Finally, the only work that addresses data heterogeneity, but not CF, in SFL (SplitFedv2) is liao2023mergesfl which merges features and adjusts the batch size.
@@ 3 Exploring CF in SFL
The goal is to train a (deep) classification model whose set of labels (classes) is $\mathcal{L}$, where $|\mathcal{L}|=L$. The training with a set $\mathcal{C}$ of clients follows the SFL’s workflow described in Sec. 1. Further, clients’ data can be seen as “tasks” (in the parlance of CL) with possibly (totally) overlapping labels.
@@ 3.1 Methodology of experiments
Models, Datasets and Cut Layer in SFL. Three ML models are employed: MobileNetV1 howard2017mobilenets, known for its efficiency on mobile and resource-constrained devices, ResNet18, and ResNet101 he2016deep. We note that ResNet101 is renowned for its depth and robust feature extraction capabilities. These models are trained on SVHN netzer2011reading, CIFAR-10 ($L=10$), CIFAR-100 krizhevsky2009learning ($L=20$ or $100$ – unless otherwise specified, we use superclasses as in ramasesh2020anatomy), and TinyImageNet le2015tiny ($L=200$). The cut layer determines the size of part-1 and part-2. A “shallow” cut layer implies a small part-1, i.e., the server trains the largest portion of the model, while a “deep” one implies a small part-2.
Based on the discussion in Sec. 1 about the intraCF, we conjecture, and later we show it, that a shallow cut better captures the impact of intraCF on the forgetting observed in SFL. For this reason, unless otherwise specified, we choose a shallow cut layer, e.g., at layer 4 (out of 26 total) in MobileNetV1, and layer 2 (out of 35 total) in ResNet101, as is common in related work thapa2022splitfed. We stress, however, that we also perform a sensitivity analysis with respect to the choice of the cut layer, while following the bottleneck approach kang2017neurosurgeon. Finally, all experiments are repeated at least $10$ times for $100$ training rounds (with varying random seed) and, then, for each metric (e.g., accuracy), the median of the observed values from all repeated experiments is computed. This is to ensure that our results are robust against data outliers leys2013detecting.
Data Partitioning and Number of Clients. We use three common data partitioning techniques for data heterogeneity. First, we employ the Dominant Label (DL) ratio method, an instance of overlapping label distribution techniques solans2024non. This uses the parameter $p\in[0,100]$ to control the percentage (i.e., $p\%$) of samples with a dominant label at each client, while the remaining samples (i.e., $(100-p)\%$) are distributed evenly among the other clients. Clearly, as $p$ increases, greater heterogeneity is produced. Throughout the paper, the notation $p\%$-DL is used, and the parameter $\phi\in\mathbb{N}$ describes the number of clients that, for each label, have this label as dominant. Hence, the total number of clients participating in the training is $C=|\mathcal{C}|=\phi\cdot L$. We also employ the Sharding and Dirichlet methods.
Metrics for Catastrophic Forgetting in SFL. A common metric of forgetting in CL is the backward transfer (BW), which has been adapted for FL as follows lee2022preservation.
@@
The quantity above measures the maximum difference of performance between each label and the best-performing one, averaged over all labels. Essentially, $\mathcal{PG}(r)$ captures the average “gap” between the per-label accuracies at round $r$.
We evaluate CF with the two metrics above, while, in Sec. 5, we employ an additional metric, the per-position accuracy. In practice, position-$k$ accuracy, for $k\in\{1\ldotp\ldotp L\}$, is the accuracy of the labels that are dominant in clients that are processed at position $k$ at the server.
@@ 3.2 Insights into Catastrophic Forgetting in SFL
Table 1: Global accuracy, Performance Gap, and Backward Transfer (reported median of the last five rounds across all runs) for cyclic & random order ($\phi$ is the scale of the clients in cyclic order). The parenthesis under the model type shows the PG and the global accuracy for IID. The upward/downward arrows indicate that large and small values are desirable, respectively.
DL Global Accuracy ($\uparrow$) Performance Gap ($\downarrow$) Backward Transfer (BW) ($\downarrow$)
@@
$\bullet$ No Specific Order (random). It is often implied in the literature that the server processes the data received by each client in a random order or a first-come-first-served (FCFS) fashion thapa2022splitfed, where the arrival time may depend on network conditions tirana2024workflow. To this end, we simulate the case of FCFS with a random function at each round.
@@
$\bullet$ Cyclic Order (in DL partitioning). This ordering guarantees structure in the way the server processes the input of different clients (at part-2). Similar in spirit to swartworth2023nearly, at each round, the server processes the clients’ input in a cycle of length $L$ based on their dominant label, i.e., first all clients whose dominant label is $X\in\mathcal{L}$ ($\phi$ in total, see Sec. 3.1), then label $Y\in\mathcal{L}$, and so on. This sequence remains the same across all rounds. In our experiments, to ensure that the results are independent of the type of label, the order (sequence) of the labels is randomly selected in each experiment.
@@
Scale of clients ($\phi$). In the cyclic order, larger values of $\phi$ imply that the data of more clients of the same type (e.g., with the same dominant label in DL partition) are processed one after the other, potentially improving global accuracy. In other words, $\phi$ gives the scale of repetition of clients of the same type. In the random order, while such repetitions may occur, they are not consistent and may vary from round to round, thus, $\phi$ is meaningful only in cyclic order.
@@
Table 1 shows the global accuracy, PG, and BW for the random and cyclic order with different values of $\phi$ in a variety of models and datasets. We observe that, in most of the lines, the best accuracy, PG, and BW are achieved by the cyclic order (for any $\phi$), with a few corner cases where random has slightly better accuracy than cyclic, but higher PG. Also, we observe that among the different values of $\phi$, lower values yield the best accuracy, indicating that diversity in the training sequence is crucial. This finding aligns with other studies in the literature on CL, e.g., lin2023theory, but have not been showcased in SFL, until now. When $\phi$ is large, the server consecutively processes $\phi$ clients’ data that have the same dominant label. This leads to more frequent updates related to a single label, allowing the model to learn this label better, but at the cost of forgetting previous ones (see Fig. 2). [^2] A small exception is observed in SVHN, which appears to be less sensitive to $\phi$ because its label distribution is uneven; thus, labels with more samples dominate model updates regardless of $\phi$. Even though training seems to be more stable (i.e., lower BW), this leads to an increased PG, which implies higher disparity among labels’ performance (leading to intraCF).
@@
Both the performance gap and the backward transfer metrics show that structured processing order policies (at the server), such as the cyclic order, lead to lower forgetting and better training stability when compared to unstructured ones, e.g., random order. This aligns with existing theoretical results on CL and linear regression evron2022catastrophic. Moreover, as we can observe already in Table 1 (in terms of accuracy and PG), many characteristics of forgetting in CL are also manifested in SFL (e.g., that larger and more complex models are more sensitive to CF ramasesh2021effect). Hence, despite the differences in the two settings (CL and SFL), forgetting in these settings seem to carry similarities, which can help to better understand the phenomenon of CF and intraCF in SFL. In the remainder of this section, we mainly focus on cyclic order with $\phi=10$, since it appears to be the most challenging case of cyclic order, in terms of accuracy and PG.
Figure 3: Global accuracy (y-axis) and PG (x-axis) for SFL with different cuts (44, $15$, and $23$, resp.) and FL/SplitFedV1.
@@
Cut Layer and IntraCF. Fig. 3 depicts the global accuracy and the PG for different cut layers (in SFL), and for SplitFedv1 thapa2022splitfed, a version of SL. In SplitFedv1, part-1 is trained as in SFL, and part-2 is trained at the server using a copy for each client, while all copies are aggregated at the end of each round. It has been shown that its performance is the same as that of FL gao2021evaluation; thapa2022splitfed, and hence, we also denote it by FL. As we go from shallow to deep cut, we notice that, in the case of high data heterogeneity, i.e., $80\%$-DL, the accuracy increases and the performance gap drops. We remind that the deeper the cut layer is, the larger part-1 becomes, i.e., more layers are trained at the clients (in parallel) and aggregated at the end of the round. Hence, a deep cut highlights the effect of CF of part-1, while a shallow cut highlights the effect of intraCF on CF. The former is related to the aggregation under data heterogeneity, while the latter to the processing order at the server. Hence, the results imply that the impact of CF in part-1 (model drifting) is smaller than the effect of CF in part-2  intraCF), i.e., the cut layer balances the impact of intraCF of part-2 and model drifting.
@@
Due to this observation, we expect that the deeper the cut layer is in SFL, the more similar the model performs to FL. We see that in terms of accuracy, SFL performs better than FL under IID data. As data heterogeneity increases, however, FL performs better ($54\%$ vs $43\%$, under $80\%$-DL), as also observed by gao2020end, as a result of aggregating the entire model. In terms of CF, we notice that, under $80\%$-DL, there is a significant performance gap between FL and SFL, for any cut layer. However, we see something unexpected: the PG of SFL with shallow cut (which showcases intraCF) is very close to that of deep cut (which showcases the CF of part-1). This means that although FL performs better under high heterogeneity, implying that aggregation has a positive impact on CF, the fact that the last layers of the model are processed at the server (sequentially) leads to high CF in SFL. The importance of the last layers of the model to the training and CF has also been shown in CL in ramasesh2020anatomy; zilly2021plasticity.

### v2

@@ Abstract
In Split Federated Learning (SFL), the clients collaboratively train a model with the help of a server by splitting the model into two parts. Part-1 is trained locally at each client and aggregated by the aggregator at the end of each round. Part-2 is trained at a server that sequentially processes the intermediate activations received from each client. We study the phenomenon of catastrophic forgetting (CF) in SFL in the presence of data heterogeneity. In detail, due to the nature of SFL, local updates of part-1 may drift away from global optima, while part-2 is sensitive to the processing sequence, similar to forgetting in continual learning (CL). Specifically, we observe that the trained model performs better in classes (labels) seen at the end of the sequence. We investigate this phenomenon with emphasis on key aspects of SFL, such as the processing order at the server and the cut layer. Based on our findings, we propose Hydra, a novel mitigation method inspired by multi-head neural networks and adapted for the SFL setting. Extensive numerical evaluations show that Hydra outperforms baselines and methods from the literature.
@@ 1 Introduction
Split Learning (SL) is a distributed learning method where part of the clients’ training is offloaded to a server (Vepakomma et al., 2018). This is particularly useful in cases where clients’ resources are insufficient to perform on-device training. In a nutshell, in SL, the deep neural network (NN) model is split into two parts of consecutive layers: part-1 and part-2, where the last layer of part-1 is called the cut layer. These parts are trained by the client and the server, respectively. There is a plethora of proposed variations for SL that mostly differ in the way the training at the server takes place. In this work, we focus on Split Federated Learning (SFL) and, in particular, SplitFedv2 (Thapa et al., 2022), which is one of the most prevailing versions (Hafi et al., 2024). In fact, thanks to its workflow, SFL has been shown to deliver a faster convergence than other variants (Thapa et al., 2022).
@@
Motivation. Part-1 is trained as in Federated Learning (FL), while part-2 is trained over the activations that clients send. This dual aspect of SFL makes it particularly susceptible to catastrophic forgetting (CF) in the presence of data heterogeneity, a common phenomenon in training NNs (Goodfellow et al., 2013). In practice, part-1 and part-2 may suffer from CF differently, as depicted in Fig. 1 (bottom). Like in FL, the local updates of part-1 drift away from the global optima as a result of the aggregation (e.g., FedAvg (McMahan et al., 2017)) at the end of every round. On the other hand, our experiments reveal a new finding: the processing order at the server has a significant impact on CF in part-2. This CF resembles the one of Continual Learning (CL) (Wang et al., 2024): the model parameters perform well for the dominant label of the last processed client, but forget the previously seen labels, i.e., in every update step at the server, part-2 tends to forget the knowledge learned during the previous steps, leading to a performance gap among labels.
@@
Fig. 2 showcases the CF in SFL as a result of the processing order at the server under non-independently and identically distributed (non-IID) data. In particular, we assume that the data of each client contains a highly represented (dominant) label. Each line depicts the accuracy of the labels that are dominant in clients that are processed at a certain position (1 to 10, in CIFAR-10 with 10 labels), where details on the experimental setup will be given in Sec. 3.1. We see that, under non-IID data, the labels that are dominant at the clients who are processed last at the server consistently outperform the other labels. The disparity in performance between the labels seen at the end of the sequence and those seen earlier is related to the CF in part-2. We identify this as intra-round catastrophic forgetting (intraCF), as it occurs at a different granularity (i.e., within a round) than the model drifting of part-1, which also contributes to the phenomenon. Moreover, the global accuracy is heavily affected, i.e., $43\%$ after 100 rounds, whereas, under IID data, the global accuracy reaches $80\%$ after 100 rounds.
@@
(i) Metrics of forgetting & Mitigation Methods. Related work on forgetting in SL, e.g., (Feng et al., 2023; Xia et al., 2025) focuses solely on the global accuracy, ignoring the disparity (gap) of performance among labels. Further, methods proposed to mitigate forgetting and/or data heterogeneity in CL and FL, such as EWC (Kirkpatrick et al., 2017) and Scaffold (Karimireddy et al., 2020), cannot be applied in the setting of SFL as the former is based on the notion of temporal data and the latter requires the entire model.
(ii) The roles of the cut layer and processing order. The majority of related work assumes that the server processes the clients’ data in a random order (Thapa et al., 2022; Liao et al., 2024) and often ignores the role of the cut layer in CF. Moreover, existing theoretical analyses on SFL are limited to convergence analysis under standard assumptions without focusing on these key elements of SFL (Han et al., 2024).
(iii) Empirical evaluation. The existing theoretical findings are often showcased in limited evaluation scenarios (e.g., partition methods), see (Li and Lyu, 2023; Han et al., 2024). Further, it is unclear whether some of the theoretical claims proved in the settings of CL and FL may hold in SFL.
@@
2. Providing valuable insights into this phenomenon (e.g., w.r.t. the impact of cut layer and processing order at the server) and comparing them with existing theoretical findings. To this end, we focus on the accuracy, the performance gap among labels, and the metric of backward transfer.
3. Proposing Hydra, a novel mitigation method that trains multiple versions of the last layers of part-2 after grouping the clients’ input based on their data distributions. Hydra is designed based on the insights from our analysis and induces minimal overhead (which can be further reduced).
@@ 2 Related Work
In this section, we discuss related work on data heterogeneity and CF in FL and SL. In FL under non-IID data, several works study CF as a result of model drifting and client selection policies. Mitigation methods utilize knowledge distillation (Lee et al., 2022; Ma et al., 2022) and regularization (Xu et al., 2022), among others. We notice, however, that most of the methods for FL require control of the entire model, e.g., (Zhao et al., 2018; Karimireddy et al., 2020; Hu et al., 2024), which is incompatible with the SFL workflow. Moreover, optimization techniques in FL (Li et al., 2020; Wang et al., 2020) assume a symmetric training workflow with full local steps or per-client regularization terms, which is also not applicable for SFL. On the other hand, replay-based methods for FL, e.g., (Qi et al., 2023; Li et al., 2024) induce computational and memory overhead. Concerning SL (in its various definitions), CF has been mostly studied in terms of global accuracy under non-IID data. Proposed solutions are based on generated data and knowledge transfer for SplitNN (Feng et al., 2023, 2024), as well as on knowledge replay for SplitFedv1 (Xia et al., 2025). Further, (Madaan et al., 2022) provides a preliminary study of the impact of the processing order in different variants of SL. Finally, the only work that addresses data heterogeneity but not CF in SFL (SplitFedv2) is (Liao et al., 2024), which merges features and adjusts the batch size.
@@ 3 Exploring CF in SFL
The goal is to train a (deep) classification model whose set of labels (classes) is $\mathcal{L}$, where $|\mathcal{L}|=L$. The training with a set $\mathcal{C}$ of clients follows the SFL workflow described in Sec. 1. Further, clients’ data can be seen as “tasks” (in the parlance of CL) with possibly (totally) overlapping labels.
@@ 3.1 Methodology of Experiments
Models, Datasets, and Cut Layer in SFL. Two ML models are employed: MobileNetV1 (Howard et al., 2017), known for its efficiency on mobile and resource-constrained devices, and ResNet101 (He et al., 2016), which is renowned for its depth and robust feature extraction capabilities. These models are trained on SVHN (Netzer et al., 2011), CIFAR-10 ($L=10$), CIFAR-100 (Krizhevsky et al., 2009) ($L=20$ or $100$ – unless otherwise specified, we use $20$ superclasses as in (Ramasesh et al., 2021a)), and TinyImageNet (Le and Yang, 2015) ($L=200$). The cut layer determines the sizes of part-1 and part-2. A shallow cut layer implies a small part-1, i.e., the server trains the model’s largest portion, while a deep one implies a small part-2.
Based on the discussion in Sec. 1 about the intraCF, unless otherwise specified, we choose a shallow cut layer, e.g., at layer 4 (out of 26 in total) in MobileNetV1, and layer 2 (out of 35 in total) in ResNet101, as is common in related work (Thapa et al., 2022). In fact, in Sec. 3, we show that a shallow cut better captures the impact of intraCF on the forgetting observed in SFL. We stress, however, that we also perform a sensitivity analysis with respect to the choice of the cut layer, while following the bottleneck approach (Kang et al., 2017). Finally, all the experiments are repeated at least $10$ times for $100$ training rounds (with varying random seeds) and, then, for each metric (e.g., global accuracy), the median of the observed values from all the repeated experiments is computed. This is to ensure that our results are robust against outliers (Leys et al., 2013).
Data Partitioning and Number of Clients. We use three common data partitioning techniques for data heterogeneity. First, we employ the Dominant Label (DL) ratio method, an instance of overlapping label distribution techniques (Jimenez G. et al., 2024). This uses the parameter $p\in[0,100]$ to control the percentage (i.e., $p\%$) of samples with a dominant label at each client, while the remaining samples (i.e., $(100-p)\%$) are distributed evenly among the other clients. Clearly, as $p$ increases, greater heterogeneity is produced. Throughout the paper, the notation $p\%$-DL is employed, and we use the parameter $\phi\in\mathbb{N}$ to describe the number of clients that, for each label, have this label as dominant. Hence, the total number of clients participating in the training is $C=|\mathcal{C}|=\phi\cdot L$. We also employ the Sharding and Dirichlet data partitioning methods.
Metrics for Catastrophic Forgetting in SFL. A common metric of forgetting in CL is the backward transfer (BW), which has been adapted for FL as follows (Lee et al., 2022):
@@
The quantity above measures the maximum difference in performance between each label and the best-performing one, averaged over all labels. Essentially, $\mathcal{PG}(r)$ captures the average “gap” between the per-label accuracies at round $r$.
We evaluate CF with the two metrics above, while, in Sec. 5, we employ an additional metric in the case of cyclic order and DL partitioning, the per-position accuracy. In practice, position-$k$ accuracy, for $k\in\{1\ldotp\ldotp L\}$, is the accuracy of the labels that are dominant in clients that are processed at position $k$ (in the cycle of length $L$) at the server.
@@ 3.2 Insights into Catastrophic Forgetting in SFL
DL Global Accuracy ($\uparrow$) Performance Gap (PG) ($\downarrow$) Backward Transfer (BW) ($\downarrow$)
@@
Table 1: Global accuracy, Performance Gap, and Backward Transfer (reported median of the last five rounds across all runs) for cyclic & random order ($\phi$ is the scale of the clients in the cyclic order). The parenthesis under the model type shows the global accuracy and the PG for IID data. The upward/downward arrows indicate that large and small values are desirable, respectively.
@@
$\bullet$ No Specific Order (random). It is often implied in the literature that the server processes the data received by each client in a random or a first-come-first-served (FCFS) order (Thapa et al., 2022), as the arrival time may depend on network conditions (Tirana et al., 2024). To this end, we simulate the case of FCFS with a random function at each round.
@@
$\bullet$ Cyclic Order. This order guarantees structure in the way the server processes the input of different clients (at part-2). Similar to (Swartworth et al., 2023), at each round, the server processes the clients’ input in a cycle of length $L$ based on the label that is highly represented in their data, i.e., first all $\phi$ clients whose highly represented label is $X\in\mathcal{L}$, then label $Y\in\mathcal{L}$, and so on. This sequence remains the same across all rounds. In our experiments, to ensure that the results are independent of the type of label, the order (sequence) of the labels is randomly selected in each experiment.
@@
Scale of clients ($\phi$). In the cyclic order, larger values of $\phi$ imply that the data of more clients of the same type (e.g., with the same dominant label in DL partition) are processed one after the other. In other words, $\phi$ gives the scale of repetition of clients of the same type. In the random order, while such repetitions may occur, they may vary from round to round, and thus, $\phi$ is meaningful only in the cyclic order.
@@
Table 1 shows the global accuracy, PG, and BW for the random and cyclic order with different values of $\phi$ in a variety of models and datasets. We observe that, in most of the lines, the best accuracy, PG, and BW are achieved by the cyclic order (for any $\phi$), with a few corner cases where random has slightly better accuracy than cyclic, but higher PG. Also, we observe that among the different values of $\phi$, lower values yield the best accuracy, indicating that diversity in the training sequence is crucial. This finding aligns with other studies in the literature on CL, e.g., (Lin et al., 2023), but has not been showcased in SFL, until now. When $\phi$ is large, the server consecutively processes $\phi$ clients’ data that have the same dominant label. This leads to more frequent updates related to a single label, allowing the model to learn this label better, but at the cost of forgetting previous ones. [^2] A small exception is observed in SVHN, which appears to be less sensitive to $\phi$ because its label distribution is uneven; thus, labels with more samples dominate model updates regardless of $\phi$. Even though training seems to be more stable (i.e., lower BW), this leads to an increased PG, which implies higher disparity among labels’ performance (leading to intraCF).
@@
Both the PG and the BW metrics show that structured processing order policies (at the server), such as the cyclic order, lead to lower forgetting and better training stability when compared to unstructured ones, e.g., random order. This aligns with existing theoretical results on CL and linear regression (Evron et al., 2022). Moreover, as we can already observe in Table 1 (in terms of accuracy and PG), many characteristics of forgetting in CL also occur in SFL (e.g., that larger and more complex models are more sensitive to CF (Ramasesh et al., 2021b)). Hence, despite the differences in the two settings (CL and SFL), forgetting in these settings carries similarities, which can help to better understand CF and intraCF in SFL.
@@
Cut Layer and IntraCF. Fig. 3 depicts the global accuracy and the PG for different cut layers (in SFL), and for SplitFedv1 (Thapa et al., 2022), a variant of SL. In SplitFedv1, part-1 is trained as in SFL, and part-2 is trained at the server using a copy for each client, while all copies are aggregated at the end of each round. It has been shown that its performance is the same as that of FL (Gao et al., 2021; Thapa et al., 2022), and hence, we also denote it by FL. For high data heterogeneity, i.e., $80\%$-DL, as we go from shallow to deep cut layer, the accuracy increases and the performance gap drops. We recall that the deeper the cut layer is, the larger part-1 becomes, i.e., more layers are trained at the clients (in parallel) and aggregated at the end of the round. Hence, a deep cut layer highlights the effect of CF in part-1, while a shallow cut layer highlights the effect of intraCF on CF. The former is related to the aggregation under data heterogeneity, while the latter to the processing order at the server. These results imply that the impact of CF in part-1 (model drifting) is smaller than the effect of CF in part-2 (intraCF), i.e., the cut layer balances the impact of model drifting of part-1 and intraCF of part-2.
@@
Due to the previous observation, we expect that the deeper the cut layer is in SFL, the more similar the model performs to FL. We see that, in terms of accuracy, SFL performs better than FL under IID data. As data heterogeneity increases, however, FL performs better ($54\%$ vs $43\%$, under $80\%$-DL), as also observed by (Gao et al., 2020), as a result of aggregating the entire model. In terms of CF, under $80\%$-DL, there is a significant performance gap between FL and SFL, for any cut layer. However, we see something unexpected: the PG of SFL with shallow cut layer (which showcases intraCF) is very close to that of deep cut layer (which showcases the CF of part-1). This means that although FL performs better under high heterogeneity, implying that aggregation has a positive impact on CF, the fact that the last layers of the model are processed at the server (sequentially) leads to high CF in SFL. The importance of the last layers of the model to the training and CF has also been shown in CL in (Ramasesh et al., 2021a; Zilly et al., 2021).
Figure 3: Accuracy (y-axis) and PG (x-axis) achieved by MobileNet/CIFAR-10 for SFL with different cuts (44, $15$, and $23$, resp.) and cyclic order ($\phi=10$), and FL/SplitFedV1.